# JEL: C81, O32

### v2

@@ A.2.1. NLM Tags
These categories are chosen to include all pubtypes with the potential to contain a publication satisfying the restriction enumerated in 3.1. In particular, we follow Feldman et al., (2019), who use multiple pubtype fields to retrieve a sample of records. Here, two authors reviewed the list of potential pubtypes to identify categories likely to include records of interest. Table A.1 reports the frequency of each NLM tag across all records in PubMed and subset to those published after 1 January 2010.
@@
Notes: Table A.1 reports the frequency and percentage of records indexed in PubMed that have been categorized by the NLM as falling into each of 18 categories. These categories are selected for their potential to contain a publication satisfying the restriction enumerated in 3.1.
@@ A.2.5. Novel Census
We add columns to Tables A.2, A.3, A.1, and A.4 that document the frequency of each flag in the data constructed for this paper (we use the conservative sample). Several facts are worth noting. First, Table A.2 suggests that—between 2010 and 2022—less than 25 percent of records identified as reporting the results of a clinical trial reported a registry identifier in their abstract. Although, in principle, registry identifiers may be reported in publication full-texts, rather than abstracts, the low frequency of identifiers in our sample suggests non-compliance with requirements, consistent with patterns in ClinicalTrials.gov documented in DeVito et al., (2020). Second, Table A.1 sheds light on the differences in trial composition that we observe in Figure 3.1: NLM tags that, ostensibly, capture clinical trials have little overlap with one another and do not fully capture the records in our sample. Of the roughly 150,000 records in our census, approximately 100,000 could be identified using NLM tags for “clinical trial” (and its variants) and “randomized controlled trial.” Observe, however, that use of such flags captures a much larger number of records than one would intend to include. Finally, Table A.4 suggests that although various text-based flags that may indicate a record is a clinical trial, in the sense of 3.1, are correlated, there is less overlap than one might expect.
@@ Appendix C Prompt Design, Fine-Tuning, and Performance Assessment
In this appendix, we detail the construction of the sample of clinical trials studied in Section 4. We begin in Section C.1 by describing the process we used to hand-label a sample of publications according to whether they satisfy the restrictions enumerated in 3.1. In Section C.2, we document poor performance of several approaches for classifying publications according to whether they belong in our sample based on standard machine learning methods. Section C.3 overviews the process of constructing prompts that perform well when used to query proprietary language models. Details associated with fine-tuning an open-source language model are given in Section C.4. Additional measurements of the performance of our model are given in Section C.5.
@@ C.1. Hand-Labeling
We devise an interface for abstract labeling, which allows us to review records quickly. Figure C.1 provides an example of this interface. A labeller reads in a set of publication records and is shown one PubMed identifier (PMID) and one abstract at a time. The labeller can select that the abstract satisfies our sample criteria, or else can indicate a reason that the abstract should be excluded. The buttons for exclusion mirror the inclusion and exclusion criteria in 3.1, using short-hand convenient for labellers. We exclude records that report, only, the protocol for a clinical trial without results (“protocol”), that report the results of an observational (“observational”) or retrospective (“retrospective”) study, that report the results of a study involving animals (“animal”), that do not include human subjects (“no human subjects”), [^67] In practice, “animal” and “no human subjects” catch different sets of records. We flag records that include any animal with “animal.” Many studies in our sample of candidate records occur in vitro—in test tubes or other laboratory-based settings. There, no living subjects of any kind are enrolled and the “no human subjects” button is used. that do not study a drug (“no drug”), that report the results of a specific clinical case or patient experience (“case report”), or that summarize the findings of existing work (“meta-analysis”). By our own estimates, the labelling interface increased the speed of labelling by a factor of seven. Each record was reviewed, and labelled, twice.
@@
Notes: Figure C.1 displays a screenshot of the application we developed to hand-label abstracts. Each abstract was labelled according to whether it satisfied the sample criteria enumerated in 3.1. If a publication did not satisfy our sample restrictions, we stipulated a reason. The user interface was developed with the “Meerkat” Python package available at the link: https://github.com/HazyResearch/meerkat.
@@ C.3. Prompt Design and Error Analysis
We test each of these three prompts in our 1000-record hand-labelled validation dataset, using both GPT-3.5 and GPT-4. We conduct a detailed error analysis, reported in LABEL:tab:Error-Analysis. To make this concrete, suppose that GPT-3.5, given Prompt 1.0 and an abstract that reports the results of a meta-analysis, erroneously classifies the record as clinical trial, per our definition. In LABEL:tab:Error-Analysis, we flag this error under Prompt 1 for GPT-3.5, as an error of type “meta-analysis.” Within meta-analysis, we assign the error to a sub-type. If the abstract explicitly includes terms such as “meta-analysis,” “literature search,” or “literature review,” we categorize this as an explicit error. If the abstract references that the publication is summarizing existing studies, or searching a database for records and collecting their findings, we categorize this as an implicit error. We devise error types based on our inclusion/exclusion criteria, in 3.1, and select sub-types based on common categories of errors.
@@
Notes: LABEL:tab:Error-Analysis categorizes errors documented in the first round of prompt iteration, using Prompts of sub-type 0. See Table 3.1 for details on model performance. For each model error—an instance in which a model returned a classification that deviated from hand-labelled data—we examined the associated record. We categorized errors by Type and Sub-Type. Types are drawn from the exclusion restrictions in 3.1. Sub-types describe consistent characteristics of publications that resulted in such an error. We report counts of errors, by model and prompt, of each Type and Sub-type.
C.4. Fine-tuning
This appendix gives technical details related to fine-tuning open-source large language models with noisy labels extracted from GPT-3.5 and GPT-4. We fine-tune pre-trained BERT models from two different architecture classes: (1) BigBird and (2) BioMedBERT (Gu et al.,, 2021; Zaheer et al.,, 2021). Both models us medium-scale Transformer architectures, i.e., between 100 and 300 million parameters, pre-trained with the masked-language modeling (MLM) objective (Vaswani et al.,, 2023; Devlin et al.,, 2019). The models differ in their architectural details and pre-training corpora.
BigBird uses a sparse attention mechanism to reduce the computational cost of processing long text sequences. The computational complexity of regular attention is quadratic in sequence length, while BigBird reduces this to linear complexity. This enables efficient handling of long documents up to 4,096 tokens. In contrast, standard BERT can only process 512 tokens. BigBird is useful because abstracts of clinical trials regularly exceed 512 tokens. Prior to pre-training, the model was warm-started from the RoBERTa checkpoint (Liu et al.,, 2019) and then pre-trained on the standard BERT corpus consisting of the Books (Zhu et al.,, 2015), CC-News (Guu et al.,, 2020), Stories (Trinh and Le,, 2019), and Wikipedia (Wikimedia Foundation,, 2023) datasets. We evaluate two model sizes: BigBird Base (125 million parameters) and BigBird Large (355 million parameters).
BioMedBert uses a domain-specific pretraining corpus sourced from articles on PubMed Gu et al., (2021). It uses the same model architecture as RoBERTa, but a different, domain-specific token vocabulary optimized for PubMed articles. Like BigBird, we evaluate two model sizes: BioMedBert Base (125 million parameters) and BioMedBert Large (355 million parameters).
We fine-tune the pre-trained architectures to classify abstracts according to the inclusion and exclusion criteria enumerated in 3.1. We replace the language modeling classification head with a binary classification head consisting of a single linear layer and softmax activation. We fine-tune the full model (i.e. not just the classification head) with cross-entropy loss. We use the Adam optimizer with learning rate $1\times 10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$ (determined via hyperparameter sweep). We use a maximum sequence length of $4{,}096$ tokens. If an abstract does not fit into the maximum sequence length provided by a model, the abstract is truncated. All models were trained on a single $V100$ GPU with 16GB of HBM.
C.5. Performance Assessment
In this appendix, we assess the performance of various fine-tuned language models for classifying whether a publication satisfies the restrictions enumerated in 3.1. Figure C.2 displays estimates of the receiver operating characteristic of several fine-tuned large language models. Here, we consider only the performance of the RoBERTa base model. The top and bottom panels are trained on noisy labels extracted from GPT-3.5 and GPT-4, respectively. We vary the quantity of noisy labels used to train the language model. Additionally, in the bottom panel, we display the performance of a fine-tuned model trained with the 1000 hand-labelled observations in the training dataset. The performance of each language model is measured in the testing dataset.
Figure C.2. Receiver Operating Characteristic by Data Size and Origin
Notes: Figure C.2 displays estimates of the the receiver operating characteristic of several fine-tuned language models. The models differ in terms of the source of their training data. Solid lines display the performance of models trained on noisy labels extracted from either GPT-3.5 or GPT-4. The dotted line displays the performance of a model trained on the hand-labeled data in the training dataset. Estimates of the true positive rate and false positive rate of GPT-3.5 and GPT-4 are indicated with black dots.
The fine-tuned models trained with noisy labels extracted from GPT-4 significantly outperform the models trained with noisy labels extracted from GPT-3.5 or with the hand-labels. There appears to be a threshold, where performance dramatically improves, at around 8,000 training labels. The models displayed in Figure 3.3 were each obtained by using all 64,000 training labels.
C.6. Final Model: Error Analysis
@@
The Conservative model incorrectly labels 27 papers. We inspect each of these errors. We review each abstract and aim to understand what might have generated an error. In roughly half of cases–13 of 27—there is a clear error. These clear errors include publications that explicitly report the results of observational studies (one error), that re-analyze existing data (eight errors), and that report literature reviews (one error). This set also includes clinical trials that satisfy all criteria in 3.1, except the treatment being studied is not a drug (three errors). In 14 cases, however, errors are associated with records that are difficult for human labellers to categorize, either because the content of the publication does not fit neatly into the inclusion and exclusion criteria implied by 3.1 or because the publication is written in a way that makes it difficult to determine details of the study. PubMed record 27880726 is illustrative. This publication studies the effect of chrysophanic acid (CA) on benign prostatic hyperplasia. The abstract is unclear about whether this treatment is studied in human subjects or in an animal model–though the publication title and full-text make clear that this was conducted in a sample of rats. Based on the abstract alone, however, nothing suggests that this is an animal study. It is excluded from both the conservative and moderate model-generated samples, but a human-labeller flagged it as satisfying our criteria.
@@ D.3.2. Text Searches for Meta-Analyses
Recall that, in the process of constructing the training data that were used to fine-tune our large language models, we identified many records in PubMed that were meta-analyses, literature reviews, etc., based on the contents of their text. We reviewed all hand-labeled records from this training process and identified a set of keywords and phrases that appeared most frequently. We constructed a list of phrases and keywords that consistently indicated, to a human labeler, that a record is a meta-analysis or review article. This list is reproduced in Footnote 51.
Using this list of search terms, we search the abstracts of all records in our citing sample. We flag any record with any of the keywords listed in Footnote 51 as a meta-analysis and flag any record that mentions at least two of the databases listed in Footnote 51 as well.

### v3

@@ A.2.1. NLM Tags
These categories are chosen to include all pubtypes with the potential to contain a publication satisfying the restriction enumerated in 2.1. In particular, we follow Feldman et al., (2019), who use multiple pubtype fields to retrieve a sample of records. Here, two authors reviewed the list of potential pubtypes to identify categories likely to include records of interest. Table A.1 reports the frequency of each NLM tag across all records in PubMed and subset to those published after 1 January 2010.
@@
Notes: Table A.1 reports the frequency and percentage of records indexed in PubMed that have been categorized by the NLM as falling into each of 18 categories. These categories are selected for their potential to contain a publication satisfying the restriction enumerated in 2.1.
@@ A.2.5. Novel Census
We add columns to Tables A.2, A.3, A.1, and A.4 that document the frequency of each flag in the data constructed for this paper (we use the conservative sample). Several facts are worth noting. First, Table A.2 suggests that—between 2010 and 2022—less than 25 percent of records identified as reporting the results of a clinical trial reported a registry identifier in their abstract. Although, in principle, registry identifiers may be reported in publication full-texts, rather than abstracts, the low frequency of identifiers in our sample suggests non-compliance with requirements, consistent with patterns in ClinicalTrials.gov documented in DeVito et al., (2020). Second, Table A.1 sheds light on the differences in trial composition that we observe in Figure 2.1: NLM tags that, ostensibly, capture clinical trials have little overlap with one another and do not fully capture the records in our sample. Of the roughly 150,000 records in our census, approximately 100,000 could be identified using NLM tags for “clinical trial” (and its variants) and “randomized controlled trial.” Observe, however, that use of such flags captures a much larger number of records than one would intend to include. Finally, Table A.4 suggests that although various text-based flags that may indicate a record is a clinical trial, in the sense of 2.1, are correlated, there is less overlap than one might expect.
@@ Appendix C Prompt Design, Fine-Tuning, and Performance Assessment
In this appendix, we detail the construction of the sample of clinical trials studied in Section 3. We begin in Section C.1 by describing the process we used to hand-label a sample of publications according to whether they satisfy the restrictions enumerated in 2.1. In Section C.2, we document poor performance of several approaches for classifying publications according to whether they belong in our sample based on standard machine learning methods. Section C.3 overviews the process of constructing prompts that perform well when used to query proprietary language models.
@@ C.1. Hand-Labeling
We devise an interface for abstract labeling, which allows us to review records quickly. Figure C.1 provides an example of this interface. A labeller reads in a set of publication records and is shown one PubMed identifier (PMID) and one abstract at a time. The labeller can select that the abstract satisfies our sample criteria, or else can indicate a reason that the abstract should be excluded. The buttons for exclusion mirror the inclusion and exclusion criteria in 2.1, using short-hand convenient for labellers. We exclude records that report, only, the protocol for a clinical trial without results (“protocol”), that report the results of an observational (“observational”) or retrospective (“retrospective”) study, that report the results of a study involving animals (“animal”), that do not include human subjects (“no human subjects”), [^55] In practice, “animal” and “no human subjects” catch different sets of records. We flag records that include any animal with “animal.” Many studies in our sample of candidate records occur in vitro—in test tubes or other laboratory-based settings. There, no living subjects of any kind are enrolled and the “no human subjects” button is used. that do not study a drug (“no drug”), that report the results of a specific clinical case or patient experience (“case report”), or that summarize the findings of existing work (“meta-analysis”). By our own estimates, the labelling interface increased the speed of labelling by a factor of seven. Each record was reviewed, and labelled, twice.
@@
Notes: Figure C.1 displays a screenshot of the application we developed to hand-label abstracts. Each abstract was labelled according to whether it satisfied the sample criteria enumerated in 2.1. If a publication did not satisfy our sample restrictions, we stipulated a reason. The user interface was developed with the “Meerkat” Python package available at the link: https://github.com/HazyResearch/meerkat.
@@ C.3. Prompt Design and Error Analysis
We test each of these three prompts in our 1000-record hand-labelled validation dataset, using both GPT-3.5 and GPT-4. We conduct a detailed error analysis, reported in LABEL:tab:Error-Analysis. To make this concrete, suppose that GPT-3.5, given Prompt 1.0 and an abstract that reports the results of a meta-analysis, erroneously classifies the record as clinical trial, per our definition. In LABEL:tab:Error-Analysis, we flag this error under Prompt 1 for GPT-3.5, as an error of type “meta-analysis.” Within meta-analysis, we assign the error to a sub-type. If the abstract explicitly includes terms such as “meta-analysis,” “literature search,” or “literature review,” we categorize this as an explicit error. If the abstract references that the publication is summarizing existing studies, or searching a database for records and collecting their findings, we categorize this as an implicit error. We devise error types based on our inclusion/exclusion criteria, in 2.1, and select sub-types based on common categories of errors.
@@
Notes: LABEL:tab:Error-Analysis categorizes errors documented in the first round of prompt iteration, using Prompts of sub-type 0. See Table 2.1 for details on model performance. For each model error—an instance in which a model returned a classification that deviated from hand-labelled data—we examined the associated record. We categorized errors by Type and Sub-Type. Types are drawn from the exclusion restrictions in 2.1. Sub-types describe consistent characteristics of publications that resulted in such an error. We report counts of errors, by model and prompt, of each Type and Sub-type.
C.4. Error Analysis
@@
The Conservative model incorrectly labels 27 papers. We inspect each of these errors. We review each abstract and aim to understand what might have generated an error. In roughly half of cases–13 of 27—there is a clear error. These clear errors include publications that explicitly report the results of observational studies (one error), that re-analyze existing data (eight errors), and that report literature reviews (one error). This set also includes clinical trials that satisfy all criteria in 2.1, except the treatment being studied is not a drug (three errors). In 14 cases, however, errors are associated with records that are difficult for human labellers to categorize, either because the content of the publication does not fit neatly into the inclusion and exclusion criteria implied by 2.1 or because the publication is written in a way that makes it difficult to determine details of the study. PubMed record 27880726 is illustrative. This publication studies the effect of chrysophanic acid (CA) on benign prostatic hyperplasia. The abstract is unclear about whether this treatment is studied in human subjects or in an animal model–though the publication title and full-text make clear that this was conducted in a sample of rats. Based on the abstract alone, however, nothing suggests that this is an animal study. It is excluded from both the conservative and moderate model-generated samples, but a human-labeller flagged it as satisfying our criteria.
@@ D.3.2. Text Searches for Meta-Analyses
Recall that, in the process of constructing the training data that were used to fine-tune our large language models, we identified many records in PubMed that were meta-analyses, literature reviews, etc., based on the contents of their text. We reviewed all hand-labeled records from this training process and identified a set of keywords and phrases that appeared most frequently. We constructed a list of phrases and keywords that consistently indicated, to a human labeler, that a record is a meta-analysis or review article. This list is reproduced in Footnote 38.
Using this list of search terms, we search the abstracts of all records in our citing sample. We flag any record with any of the keywords listed in Footnote 38 as a meta-analysis and flag any record that mentions at least two of the databases listed in Footnote 38 as well.